OpenStudy (anonymous):

Multiply (2.4 ⋅ 1014) ⋅ (4 ⋅ 107). Express the answer in scientific notation.

OpenStudy (brooke2353):

Okay well what is 2.4 times 1014

OpenStudy (anonymous):

Multiply the coefficient terms and then multiply the powers of 10. 2.4 * 4 = 9.6 10^14 * 10^7 = 10^(14+7) = 10^(21) So the answer is 9.6 x 10^21